OpenStudy (anonymous):

Helppppppp Find k if the lines 3x-5y=9 & 2x+ky=11 are a)parallel and b)perpendicular

OpenStudy (anonymous):

3x - 5y = 9 2x + ky = 11 Express the equations in the form of y = mx + c, where m is the gradient and c is the y-intercept. a) 3x - 5y = 9 5y = 3x - 9 y = 3x / 5 - 9 / 5 2x + ky = 11 ky = -2x + 11 y = -2x / k + 11 / k Since the lines are parallel, the gradient are the same. 3 / 5 = -2 / k k = -2 / (3 / 5) = -10 / 3 b) Now, the lines are perpendicular, which means the gradients are the negative reciprocals of each other. 3 / 5 = -1 / (-2 / k) = k / 2 k = 6 / 5